All of the following are measures of lengths. Rank the values from smallest to largest:
7.49x105 nm, 34500 inches, 0.45 miles, 885 yards.
1
Expert's answer
2019-10-02T08:39:56-0400

Converting all into metres

"(a)=\ 7.49\times 10^{5}\times 10^{-9}=7.49\times 10^{-4}"

"(b)=\ 34500\times 2.54\times 10^{-2}=876.3\ m\\(c)=\ 0.45\times 1609.344=724.2048\ m"

"(d)=\ 885\times 0.9144\ = 809.244\ m"

"So,
(a)\lt(c)\lt(d)\lt(b)"